|EGI Inspire Main page|
|Inspire reports menu:||Home •||SA1 weekly Reports •||SA1 Task QR Reports •||NGI QR Reports •||NGI QR User support Reports|
Progress of SA1 issues
(SA1) PQ11: NGI operations sustainability. From Milestones MS123 (Int task assessment), there is no indication that the sustainability of NGIs has improved since the last assessment in September 2012. This was clearly indicated in the conclusions of the milestone above.
- D4.8. Tiziana doing the internal review. Reviewing various figures for consistency with the PY2 report. I hope this document can go to external review after the AMB tomorrow
- MS425 Operational Level Agreements (OLAs) within the EGI production infrastructure. Milestone draft well advanced.
SA1.1 Activity Management
- Operations Portal: discussion of VO availability/reliability computation algorithm for implementation of the first prototype
- fixing many issues with incomplete/incorrect information in VO ID cards
- VO registration/deregistration management
- D4.8 and MS425 editing
- Collection of input for TF13 programme
- Set of task force for requirements gathering for A/R computation mini project
- Editing of material for EGI newsletter
- usual routine activities
- preparing for face to face meeting of the CSIRT team in Linkoping, Sweden (24-25 April)
- this will include participation by staff from PRACE and EUDAT
- work on reporting module for Security Service Challenges (and preparing reports)
- central emergency suspension
- 1 new vulnerability reported
- SVG monthly meeting
- New RAT member invited and agreed to join (Simon Fayer from Imperial)
- Updated compromised certificate procedure
SA1.3 Staged rollout
- Continuing of the activities regarding the UMD-2
- Preparation of a new release in order to scope a new security vulnerability reported in the blah component.
- Final preparation for the major release UMD-3 which contains products from EMI-3 topics :
- Collaboration with the CSIRT team for the elaboration of an wiki page regarding the Argus interoperability with the remaining middleware.
- wiki page can be found here
- assessment of status of VERCE needs about iRODS
- request of input on availability of a SSM2 central DB (in production) from collection of usage records from globus resources
SA1.4 Central tools
Operations as usual.
Repository Ongoing problems with the APEL consumer server experienced, which receives the daily load of records from the APEL clients. Despite expectations, the system did not "catch up" as it did in the past. The network problem which occurred over the Easter weekend was the cause of the latest unscheduled downtime.
More disk was added so that the database table can be optimized without filling up the disk, as this is where the I/O contention is. Following discussions with other members of the systems and databases teams at STFC, the service will be migrated to a physical server as it is currently running on a VM. The physical server is also a higher spec. system with SSDs so this should improve the performance issues we are seeing with the additional data load received. The installation of the new system started. The new system is planned to be rolled to production this week assuming verification is passed. The switchover will require about 4 hours downtime on the APEL broker and consumer systems to handle data migration.
- Successful testing by client site of the APEL 3 client, looking to move to production in next week.
- Other external CPU accounting clients (ARC) using SSM 2.0 require further verification of records as VO data not always correct and some missing fields, they are working on this and we will verify new records over the next week.
Portal No report received
- Ticket monitoring
- Implemented wiki pages describing work flows for tickets waiting for submitter input or PT input:
Discussion of these new proposed workflows with UCB.
- Phone conference with Savannah and JIRA admins on migration of GGUS shopping list from Savannah to JIRA
- Support on how using the report generator
- Prepared GGUS release; installed release on test instance
- Weekly shopping list phone meeting
- Internal wrap-up of EGI CF
- COD chairing of Nagios probe working group meeting
- followup of new sites with new EMI-1 alarms for software retirement
- follow up of RPI, sites to be suspended, top BDII A/R, Unknowns
|DMSU tickets flow Apr 14-20|
|back to tpm||0|
|reassigned to 3rd level||9|
|open DMSU tickets status|
|waiting for reply||4|
no report received
SA1.8 Availability and core services
- Initialization of Registration Authority for EGI.eu Catch-all CE in Nigeria
- Started testing of EMI-3 VOMS service endpoint for Dteam VO migration
- Ongoing work on MS425 document (due end of April 2013)
- ENVRI WP3 meeting, May, Amsterdam